Calculate the mass (in grams) of magnesium chloride (MgCl2) that would be needed to prepare 150 mL of a 20% by mass aqueous solution of the salt. The density of the solution is 1.1 g/mL.


1
Expert's answer
2021-02-22T06:09:49-0500

Given the density of the solution = 1.1 gram/ ml

So the mass of 150 ml of aqua solution = 150 x density of the solution

 = 150 ml X 1.1 gram /ml

=165 grams

According to the given data,

Magnesium chloride present in the solution = 20% by mass of the solution

 ="\dfrac{Mass of solution \times 20}{100}" gram

= "\dfrac{165 \times 20}{100}" gram

 =33 grams


hence, mass of the magnesium chloride needed to prepare a solution of 20% of magnesium chloride by mass = 33 grams of magnesium chloride